OpenStudy (anonymous):

[-5ln(2/3) - 2ln(5/3)] - [-5ln(5/6) - 2ln(4/3)] = ln (125/64) How? I can't get it. Thanks everyone

OpenStudy (anonymous):

Despite it looks a bit like -Sin, it's 5 ln ...

OpenStudy (anonymous):

Okay everyone, I've got it. Thanks anyhow!

OpenStudy (anonymous):

[-5ln(2/3) - 2ln(5/3)] - [-5ln(5/6) - 2ln(4/3)] = [5ln(3/2)+2ln(3/5)]-[5ln(6/5)+2ln(3/4)] =5ln((3*5)/(2*6))+2ln((3*4)/(3*5)) =5ln(5/4)+2ln(4/5)=3ln(5/4)=ln(125/64)